The maximum speed and acceleration of a simple harmonic oscillator are 0.87 m/s and 1.48 m/s^2

A. Find the amplitude of this motion.

B. Find the period of this motion.

Express your answers with the appropriate units.
Answer #1


maximum speed vmax = Aw = 0.87 ..(1)

amax = Aw^2 = 1.48 ..(2)

from (2)/(1)

w = 1.48/0.87 = 1.7 rad/s

from (1) A*1.7 = 0.87

A =0.512 m

(b) w = 2pi/T = 1.7

2*3.14/T = 1.7

T = 3.7 s